OpenStudy (anonymous):

HELPPP !!!!!!!!!!!!! (3x^3y^4)^2 Simplify

OpenStudy (jhannybean):

@SithsAndGiggles

OpenStudy (anonymous):

Use the properties of exponents: \[\left(a^b\right)^c=a^{bc}\]

OpenStudy (anonymous):

ok But I Still Need Help

OpenStudy (jhannybean):

Ohh... So you would have \(\large 3x ^{3y ^{8}}\) right?...

OpenStudy (anonymous):

I think the given expression says \[\left(3x^3y^4\right)^2\] not \[\left(3x^{3y^4}\right)^2\]@Jhannybean

OpenStudy (jhannybean):

Ahh... no parenthesis around the powers kind of threw me off.

OpenStudy (anonymous):

Even so, no, the exponent would also be distributed to every factor: \[\left(abc\right)^n=a^nb^nc^n\] So you have \[\left(3x^3y^4\right)^2=3^2\left(x^3\right)^2\left(y^4\right)^2\]